### v3

@@
The Cost of Coming Outthanks:  We especially would like to thank Michael Lechner and Franco Peracchi for feedback and suggestions. We are also grateful to Jaime Arellano-Bover, Nora Bearth, Jonathan Chassot, Caroline Coly, Daniel Goller, Eric Guan, Giorgio Gulino, Moritz Janas, Giovanni Mellace, David Neumark, Paolo Pinotti, Mounu Prem, Dario Sansone, Erik-Jan Senn, Andrea Smurra, seminar participants at University of Rome Tor Vergata and SEW-HSG research seminars, and conference participants at the 2nd Rome Ph.D. in Economics and Finance Conference, the SES 2024, the 29th IPDC, and the ESPE 2025 for comments and discussions.
@@
(August 7, 2025)
The fear of social stigma leads many individuals worldwide to hesitate in disclosing their sexual orientation. Since concealing identity is costly, it is crucial to understand the extent of anti-LGB sentiments and reactions to coming out. This paper uses an innovative data source from a popular online game together with a natural experiment to overcome existing data and endogeneity issues. We exploit exogenous variation in the identity of a character to identify the effects of coming out on players’ revealed preferences for that character across diverse regions globally. Our findings reveal a substantial and persistent negative impact of coming out.
Keywords: LGB economics, social stigma, concealable stigma.
@@ 1 Introduction
The fear of social stigma significantly shapes human behavior, steering individuals to conceal actions, behaviors, or identity [137, 163]. One important part of identity that has increasingly captured public and academic attention over the last two decades is an individual’s sexual orientation [[, see, e.g.,]]badgett2023review. Despite significant progress in advancing lesbian, gay, and bisexual (LGB) rights, pervasive anti-LGB sentiments persist in many countries, leading many individuals worldwide to hesitate in openly disclosing their sexual orientation due to lingering fears [174] and anticipated discrimination [191].
Existing research indicates that concealing one’s identity comes with a cost [143]. For instance, the stress of hiding one’s sexual orientation is one of the potential minority stressors [142, 146] that explain the higher prevalence of mental disorders among LGB individuals compared to their heterosexual counterparts [[, see, e.g.,]]pachankis2020sexual. Therefore, understanding the extent of stigma and the challenges faced by individuals upon coming out is crucial to fostering the groundwork for supportive work and living environments [180].
In this paper, we use an innovative data source from an online video game together with a natural experiment to assess the stigma attached to sexual minority status by observing individuals’ responses to sexual minority disclosure. Existing studies examining anti-LGB sentiments, which we review below, usually face at least one of two limitations. First, they rely on a selection-on-observables identification strategy that is likely to suffer from omitted variable bias. Second, they depend on survey data where individuals have the discretion to choose whether to disclose their sexual orientation, thereby introducing additional endogeneity issues [165, 201]. Our use of video game data effectively avoids these selection concerns, and the natural experiment we leverage allows for a credible identification of the responses to revealing sexual minority status.
@@
To strengthen the credibility of preferences towards LGB status as the primary explanation for the estimated effects, it is crucial to ensure that players’ decisions to switch from the character are not influenced by other factors. We address and eliminate several alternative channels, thereby enhancing the plausibility of social stigma as the primary explanation for the observed behavior. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character, thus dismissing the possibility that gameplay factors are the driving force behind the players’ observed behavior. Additionally, we demonstrate that players are not leaving the game after the disclosure but are shifting their focus to other characters. Third, we provide evidence that switching to other characters does not affect the performance of the players involved, highlighting that the decision to abandon the character is not driven by performance considerations. Fourth, we dismiss the possibility that the release of a new character after the disclosure explains the estimated effect.
@@
Finally, we exploit the presence of other playable characters with sexual minority status to show that LGBT Pride Month, which started on the day of the announcement, is unlikely to explain our findings. To do so, we introduce a theoretical framework that formalizes the existence of two “simultaneous treatments”—the disclosure of the character’s sexual orientation and the start of LGBT Pride Month. We outline sufficient assumptions that enable us to separate the impacts of these treatments on players’ preferences for the character.111 See, e.g., [197] for a discussion on “simultaneous treatments” and methodologies for disentangling their effects under a difference-in-differences identification strategy. The empirical results support the interpretation that the estimated effects are driven by the character’s disclosure.
While our setting is unconventional, it presents unique advantages for the identification of sentiments towards LGB identity. First, video games offer a controlled research environment, enabling the observation of behaviors that might be challenging to capture through traditional methods [204].222 A similar argument underlies several studies that use data from the professional sports industry to examine, for example, discrimination [150, 170], incentive effects [152], and other behavioral patterns [203]. Second, they allow us to leverage objective measures of behavior and identity, circumventing the limitations associated with self-reported identity in surveys [165, 201]. Third, online gaming platforms offer the benefit of anonymity, minimizing social desirability bias and increasing the likelihood of individuals disclosing sensitive information such as their true attitudes toward sexual minority groups.
@@
Despite these advantages, our study’s findings should be interpreted with three important caveats. First, the sample of participants in our setting is unlikely to be representative of the general population. The demographic composition of online gaming communities may differ systematically from broader societal distributions, potentially affecting the external validity of our findings. Second, our analysis focuses on responses to a male character revealing his sexual minority status. Given prior evidence of differential labor market discrimination against gay men and lesbian women [192], it is plausible that reactions would vary if a female character disclosed the same identity. Third, our study examines preferences for a fictitious character rather than an actual human being. As a result, factors such as the character’s perceived prestige or popularity may influence individuals’ responses, making our setting most analogous to the context of supporting a well-known public figure, such as a politician or professional athlete.
@@
Our paper contributes to three distinct strands of the literature. First, it relates to a growing literature studying the economics of LGBTQ+ individuals [[, see, e.g.,]]badgett2021lgbtq, badgett2023review. The current body of research primarily focuses on measuring discrimination against LGB individuals by comparing their labor market outcomes with those of non-minority individuals with similar observable characteristics, mostly documenting wage penalties for gay and bisexual men and wage premiums for lesbian women [[, see, e.g.,]]badgett1995wage, martell2021labor, carpenter2017does, carpenter2023orient.333 The only study finding a wage premium for gay men is that of [164]. [202] does not find a wage penalty but finds a lower probability of being full-time employed and a higher probability of being a victim of work-related violence for gay men. There is also evidence that transgender workers face earning and employment penalties [167, 186]. However, despite their valuable contributions, these studies suffer from the endogeneity and selection issues discussed above that hinder the ability to draw causal inferences from their findings.444 A related literature studies whether changes in laws and norms affect labor market outcomes and attitudes towards LGB individuals [[, see, e.g.,]]burn2018not, sansone2019pink, ofosu2019same, burn2020relationship, delhommer2020effect, deal2022bound, deal2023heterogeneity. [161] show that a randomized intervention that encourages actively taking the perspective of others can reduce transphobia. To tilt towards a more causal interpretation, other studies use correspondence designs to probe into hiring discrimination against LGB individuals, consistently revealing that LGB job candidates are less likely to be invited for interviews or offered job opportunities [147, 148, 153, 155, 157]. Nevertheless, a significant challenge lies in communicating that an individual belongs to a sexual orientation minority group, since such information is not typically included in job applications. This raises the question of whether the observed results are affected by the choice and nature of the signal used [162].
@@
We make several contributions to this literature. First, our use of data from an online video game allows us to overcome the endogeneity issues of previous studies and to identify preferences toward LGB status. Second, to the best of our knowledge, our study is the first to investigate the immediate reactions to coming out. Due to the substantial anecdotal and scientific evidence that individuals often hide their sexual orientation status [[, see, e.g.,]]badgett2020economic and the prevalent evidence of the associated costs [146, 177], several studies have explicitly focused on investigating the determinants and incentives of coming-out decisions.555 [183] investigate the impact of same-sex marriage legalization on coming-out decisions using a revealed preference mechanism inferred from data on Catholic priests’ vow of celibacy, finding reduced demand for priestly studies after the adoption of same-sex legalization. [188] explore spillover effects of coming-out decisions using Twitter data, discovering positive externalities, as exposure to peers coming out is associated with a higher probability of individuals coming out themselves. [191] conduct a lab experiment demonstrating that individuals strategically hide their sexual orientation in anticipation of discrimination in prosocial behavior, a result consistent with that of [189]. Our study complements these studies by focusing on responses to revealing sexual minority status, potentially providing a rationale for the observed underrepresentation of individuals with sexual minority status in many areas.
Second, our findings relate to the broader literature on discrimination. Early work by [138] paved the way for an extensive literature investigating discrimination instances, mostly based on gender and ethnicity, across diverse economic domains [[, see, e.g.,]]arnold2022measuring, kuhn2013gender. A substantial portion of the empirical evidence stems from field experiments, in which researchers use audit and correspondence studies to isolate the causal impact of identity on behavior [[, see, e.g.,]]ayres1995race, oreopoulos2011.666 For recent surveys, see [162] and [168]. [190] provides a detailed review of traditional statistical and taste-based discrimination models, along with a discussion of recent theories that expand on these models. Our study contributes by exploring sentiments against sexual orientation minorities. The concealable nature of sexual orientation allows individuals to anticipate discrimination and strategically choose to hide their identity [191]. This complicates the use of traditional methods to investigate discrimination against individuals based on their sexual orientation and necessitates innovative strategies for understanding sentiments against groups with stigmatized identities, particularly in environments prone to discrimination.
Third, our study contributes to the literature on video games and social interaction platforms [199]. The recognition of video game data’s potential for research is increasing among economists, who are already capitalizing on the abundance and quality of the available data.777 For example, [169] explore the impact of diversity on team performance using data from an online video game, and [195] investigate the effects of artificial intelligence on collaborative production dynamics in controlled laboratory settings using a team-based video game. [144] utilize a self-designed video game to study racial discrimination. However, despite its industry size and the significant fraction of time spent on playing video games, the scientific literature on these social platforms is only slowly increasing.888 [154] highlight that, according to the American Time Use Survey, almost 10 % of leisure time is spent on playing games and computer use. Global revenues of the video game industry were around \$180 billion in $2020$. [196] investigate consumer behavior in the video game industry. They examine the impact of marking products with a gay label on consumers’ demand, finding a significant, albeit short-lived, decrease in demand following the introduction of the gay label. [200] study entertainment preferences in the context of video games and test economic theories of belief-based utility using data from League of Legends. Our study complements this line of research by expanding the understanding of individuals’ behavior within these virtual environments.
@@
The rest of the paper unfolds as follows. Section 2 describes the key elements of League of Legends that are relevant to our study and outlines the natural experiment we leverage to identify the effects of coming out. Section 3 introduces the data and explains the methodology we use to isolate the effects of coming out. Section 4 presents our main results. Section 5 examines the underlying mechanisms driving the estimated effects. Section 6 concludes.
@@ 2.1 League of Legends
League of Legends is a globally prominent multiplayer online game developed and published by Riot Games. Originally launched on October $27^{th}$, $2009$, the game attracted a vast player base, averaging $180$ million monthly active players as of $2022$ and peaking at $14$ million players in a single day. It has also achieved substantial financial success, generating \$1.8\\$1.8 billion in revenue in that same year.999 See, e.g., https://prioridata.com/data/league-of-legends.
League of Legends demands significant time and cognitive engagement from its user base. In our sample alone—which follows a subset of players over a seven-month period—users collectively spent 647,966647{,}966 hours in matches. Valuing this gameplay time at the 2022 U.S. federal minimum wage of \$7.25\\$7.25 per hour implies an aggregate opportunity cost of \$4.7\\$4.7 million. This figure represents a conservative lower bound of time spent on the game, as it excludes substantial time spent on pre- and post-match activities.
In League of Legends, players are divided into two teams of five players each to compete in matches with the aim of destroying the opposing team’s base. Players in each team sort themselves into one of five roles. These roles represent crucial strategic positions, each requiring specific playstyles and contributing differently to the team’s final objective.
Before a match begins, players must select a playable character to control during the match. In our analysis, we measure players’ revealed preferences for a specific character by quantifying how frequently they select that character for their matches. Our objective is to investigate whether these preferences undergo any shifts following the disclosure of the character’s sexual orientation. Thus, we devote the rest of this section to exploring the design of characters in League of Legends and the process through which players select their characters for matches.
Each character has a unique set of skills and abilities and is specifically designed to excel in one or two of the distinct roles that players can take on within the team.101010 These roles are labeled top, jungle, mid, bottom, and support. Additionally, characters are crafted with a rich background that adds a narrative dimension to the game but does not have any impact on the game’s mechanics. This is achieved through the creation of detailed biographies and short stories that provide players with a deeper understanding of the character’s history and motivations, thus offering players the opportunity to connect with their chosen characters on a more personal level.111111 The list of all characters along with details about their abilities and histories is available at https://www.leagueoflegends.com/en-gb/champions/.

### v4

@@
Behavioral Consequences of Sexual Orientation Disclosure in a Large-Scale Digital Environmentthanks: A previous version of this paper circulated under the title The Cost of Coming Out. We especially would like to thank Michael Lechner and Franco Peracchi for feedback and suggestions. We are also grateful to Jaime Arellano-Bover, Nora Bearth, Jonathan Chassot, Caroline Coly, Daniel Goller, Eric Guan, Giorgio Gulino, Moritz Janas, Giovanni Mellace, David Neumark, Paolo Pinotti, Mounu Prem, Dario Sansone, Erik-Jan Senn, Andrea Smurra, seminar participants at University of Rome Tor Vergata and SEW-HSG research seminars, and conference participants at the 2nd Rome Ph.D. in Economics and Finance Conference, the SES 2024, the 29th IPDC, and the ESPE 2025 for comments and discussions.
@@
Many individuals hesitate to disclose their sexual orientation, anticipating that disclosure may alter how others respond to them. At the same time, concealing one’s identity can entail substantial personal and social costs. Understanding how others react to sexual orientation disclosure is therefore central to evaluating the broader consequences of coming out. This paper uses an innovative data set from a popular online video game together with a natural experiment to causally identify behavioral responses to sexual minority disclosure. We exploit exogenous variation in the identity of a playable character to identify the effects of coming out on players’ revealed preferences for that character across diverse regions globally. Our findings reveal a substantial and persistent negative impact of coming out.
Keywords: LGB economics, sexual orientation disclosure, homophobia, revealed preferences, online gaming, videogames.
@@ 1 Introduction
Sexual orientation is an important component of identity that shapes social and economic interactions [7]. However, while legal protections and social acceptance of lesbian, gay, and bisexual (LGB) individuals have expanded in many countries, negative attitudes and unequal treatment persist in many contexts [8, 6]. As a result, many individuals anticipate adverse reactions and hesitate to disclose their sexual orientation [4].
Yet, concealment can be costly [3]. The stress associated with hiding one’s sexual orientation constitutes a minority stressor [36, 37] and contributes to documented disparities in mental health outcomes between LGB and heterosexual individuals [40]. Therefore, understanding the behavioral and social responses faced by individuals upon coming out is crucial to fostering the groundwork for supportive work and living environments.
In this paper, we use an innovative data source from an online multiplayer video game together with a natural experiment to provide causal evidence on the behavioral costs of sexual orientation disclosure in an anonymous, high-frequency revealed-preference setting. Existing studies examining anti-LGB sentiments, which we review below, usually face at least one of two limitations. First, they rely on a selection-on-observables identification strategy that is likely to suffer from omitted variable bias. Second, they depend on survey data where individuals have the discretion to choose whether to disclose their sexual orientation, thereby introducing additional endogeneity issues [22, 32]. Our use of video game data effectively avoids these selection concerns, and the natural experiment we leverage allows for a credible identification of the behavioral responses to revealing sexual minority status.
@@
To strengthen the interpretation of the estimated effects as responses to sexual minority disclosure, it is crucial to ensure that players’ decisions to switch from the character are not influenced by other factors. We address and eliminate several alternative channels, thereby enhancing the plausibility that the observed behavior reflects a response to the disclosure itself rather than confounding gameplay or design factors. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character, thus mitigating the concern that gameplay factors are the driving force behind the players’ observed behavior. Additionally, we demonstrate that players are not leaving the game after the disclosure but are shifting their focus to other characters. Third, we provide evidence that switching to other characters does not affect the performance of the players involved, highlighting that the decision to abandon the character is unlikely to be driven by performance considerations. Fourth, we dismiss the possibility that the release of a new character after the disclosure explains the estimated effect.
@@
Finally, we exploit the presence of other playable characters with sexual minority status to show that LGBT Pride Month, which started on the day of the announcement, is unlikely to explain our findings. To do so, we introduce a theoretical framework that formalizes the existence of two “simultaneous treatments”—the disclosure of the character’s sexual orientation and the start of LGBT Pride Month.111 See, e.g., [47] for a discussion on “simultaneous treatments” and methodologies for disentangling their effects under a difference-in-differences identification strategy. We outline sufficient assumptions that enable us to separate the impacts of these treatments on players’ preferences for the character. The empirical results support the interpretation that the estimated effects are driven by the character’s disclosure.
The observed decline in character usage may reflect negative preferences or prejudice toward sexual minorities [12, 6], but it may also arise from reduced identification following disclosure of an identity that differs from that of many players [35]. Both mechanisms imply reduced demand conditional on disclosure, even when material payoffs remain unchanged.222 In models of identity utility [3], individuals derive utility not only from material outcomes but also from alignment with others’ identities; disclosure may therefore reduce perceived similarity and, in turn, demand, even absent animus [39]. Relatedly, homophily models of in-group favoritism [24, 10] imply that disclosure can shift a character from an implicitly majority-aligned identity to an out-group identity, lowering utility through either out-group aversion or in-group preference. As in much of the discrimination literature, reduced-form evidence cannot distinguish between these mechanisms. Our contribution is therefore to identify the causal behavioral response to identity disclosure, capturing economically meaningful shifts in revealed preferences.
While our setting is unconventional, it provides unique advantages for causally identifying behavioral responses to sexual minority disclosure. First, video games offer a highly controlled research environment in which behavior can be observed in real time, often in ways that are difficult to capture using traditional data sources [41].333 A similar argument underlies several studies that use data from the professional sports industry to examine, for example, discrimination [46, 44], incentive effects [45], and other behavioral patterns [18]. Second, they allow us to leverage objective measures of behavior and identity, unlike survey-based studies, which rely on self-reported identity and are vulnerable to reporting incentives and social desirability biases [22, 32]. Third, our setting provides data on high-frequency longitudinal revealed preferences over an economically meaningful time horizon. We can track character selection decisions at a daily frequency, enabling a granular assessment of both the immediate and persistent behavioral responses to disclosure. Such temporal resolution is rarely available in traditional social or labor market environments. Fourth, online gaming platforms offer the benefit of anonymity, minimizing social desirability bias and increasing the likelihood that observed behavior reflects individuals’ genuine choices rather than socially desirable responses.444 Moreover, a growing body of work documents prejudice, harassment, and exclusionary behaviors in online gaming communities. Surveys and observational studies show that hostile, sexist, and sexually prejudiced behaviors are common in online multiplayer games, with negative comments about sexual orientation frequently observed [51, 30, 34]. This evidence motivates online gaming environments as particularly salient settings for studying behavioral responses to sexual minority disclosure in a setting characterized by large-scale interaction and anonymity.
@@
Despite these advantages, our study’s findings should be interpreted with three important caveats. First, the sample of participants in our setting is unlikely to be representative of the general population. Our data are fully anonymized and do not contain individual demographic information. However, external evidence consistently indicates that the League of Legends player base is predominantly male (around $80\%$) and relatively young, with the majority of players between 16 and 34 years old [53, 28]. Our estimates therefore most plausibly reflect behavioral responses within a population of young male players rather than the broader population. At the same time, this group is typically underrepresented in traditional online and telephone surveys. Observing revealed behavior in large-scale digital data thus provides a complementary perspective on social preferences within a demographic that is otherwise difficult to study using standard survey instruments. Second, our analysis focuses on responses to a male character revealing his sexual minority status. Given prior evidence of differential labor market discrimination against gay men and lesbian women [6], it is plausible that reactions would vary if a female character disclosed the same identity. Third, our study examines preferences for a fictitious character rather than an actual human being. As a result, factors such as the character’s perceived prestige or popularity may influence individuals’ responses, making our setting most analogous to the context of supporting a well-known public figure, such as a politician or professional athlete.555 In male-dominated domains such as professional sports, public disclosure of sexual orientation remains relatively rare, consistent with the view that identity revelation in such contexts may entail substantial social or reputational costs.
@@
Our paper contributes to three distinct strands of the literature. First, it relates to a growing literature studying the economics of LGBTQ+ individuals [[, see, e.g.,]]badgett2021lgbtq, badgett2023review. The current body of research primarily focuses on measuring discrimination against LGB individuals by comparing their labor market outcomes with those of non-minority individuals with similar observable characteristics, mostly documenting wage penalties for gay and bisexual men and wage premiums for lesbian women [[, see, e.g.,]]badgett1995wage, martell2021labor, carpenter2017does, carpenter2023orient.666 The only study finding a wage premium for gay men is that of [20]. [50] does not find a wage penalty but finds a lower probability of being full-time employed and a higher probability of being a victim of work-related violence for gay men. There is also evidence that transgender workers face earning and employment penalties [29, 21]. However, despite their valuable contributions, these studies suffer from the endogeneity and selection issues discussed above that hinder the ability to draw causal inferences from their findings.777 A related literature studies whether changes in laws and norms affect labor market outcomes and attitudes towards LGB individuals [[, see, e.g.,]]burn2018not, sansone2019pink, ofosu2019same, burn2020relationship, delhommer2020effect, deal2022bound, deal2023heterogeneity. [17] show that a randomized intervention that encourages actively taking the perspective of others can reduce transphobia. To tilt towards a more causal interpretation, other studies use correspondence designs to probe into hiring discrimination against LGB individuals, consistently revealing that LGB job candidates are less likely to be invited for interviews or offered job opportunities [54, 26, 52, 2, 27]. Nevertheless, a significant challenge lies in communicating that an individual belongs to a sexual orientation minority group, since such information is not typically included in job applications. This raises the question of whether the observed results are affected by the choice and nature of the signal used [11].
@@
We make several contributions to this literature. First, our use of data from an online video game allows us to overcome the endogeneity issues of previous studies and to identify behavioral responses to sexual minority disclosure. Second, to the best of our knowledge, our study is the first to investigate the immediate reactions to coming out. Due to the substantial anecdotal and scientific evidence that individuals often conceal their sexual orientation status [[, see, e.g.,]]badgett2020economic and the prevalent evidence of the associated costs [37, 40], several studies have explicitly focused on investigating the determinants and incentives of coming-out decisions.888 [49] investigate the impact of same-sex marriage legalization on coming-out decisions using a revealed preference mechanism inferred from data on Catholic priests’ vow of celibacy. [31] explore spillover effects of coming-out decisions using Twitter data, discovering positive externalities, as exposure to peers coming out is associated with a higher probability of individuals coming out themselves. [4] conduct a lab experiment demonstrating that individuals strategically hide their sexual orientation in anticipation of discrimination in prosocial behavior, a result consistent with that of [33]. Our paper complements these studies by focusing on behavioral responses to sexual minority disclosure, thereby shedding light on how disclosure may affect subsequent social and economic interactions.
Second, our findings relate to the broader literature on identity-based behavior. Building on the seminal framework of [9], a large empirical literature, primarily focused on race and gender, documents differential treatment across labor and other markets [[, see, e.g.,]]arnold2022measuring, kuhn2013gender. Much of this evidence relies on audit and correspondence designs that exploit observable identity characteristics to isolate causal effects [[, see, e.g.,]]ayres1995race, oreopoulos2011. Sexual orientation, however, differs from most commonly studied traits in that it is often concealable. Individuals may anticipate discrimination and strategically withhold disclosure [4], rendering identity endogenous and complicating standard empirical approaches. When identity is revealed, both discrimination models and homophily frameworks predict sorting and demand shifts along group lines, albeit through distinct mechanisms. Our setting leverages an exogenous disclosure shock in a context where identity would otherwise remain hidden, thereby allowing us to identify the net behavioral response to sexual orientation revelation. In doing so, we provide causal evidence on how identity disclosure reshapes demand.
Third, our study contributes to the emerging literature on social behavior in online gaming environments by providing causal evidence on how sexual orientation disclosure affects revealed behavior. Despite the economic size of the industry and the substantial fraction of time devoted to video games, the scientific literature on social interactions within these platforms is only slowly increasing.999 [1] report that, according to the American Time Use Survey, almost $10\%$ of leisure time is spent on gaming and computer use. Global video game revenues exceeded $\mathdollar 180$ billion in $2020$. Existing work documents harassment, exclusion, and prejudice in multiplayer games [[, see, e.g.,]]tang2020investigating, gillin2024attitudes, marques2024positive, but causal evidence on behavioral responses in these environments remains scarce. Our setting allows us to exploit a plausibly exogenous sexual orientation shock and trace its impact on actual choice behavior at scale, providing novel empirical evidence on how individuals adjust their behavior following sexual orientation disclosure in digital communities.
Our paper further complements the growing literature that uses video games as empirical laboratories for economic and social behavior. Early work established the credibility of gaming environments as sources of meaningful behavioral data and as part of broader patterns of social and cultural participation, showing that behavior in virtual environments reflects stable underlying preferences rather than purely recreational noise [14, 15, 16]. Building on this foundation, recent studies exploit large-scale game data to study team performance and diversity [42], human–AI collaboration [25], racial discrimination [23], consumer responses to identity labeling [43], and belief-based utility theories [28].101010 [43] are most closely related to our study. However, our institutional setting, data and design allow us to causally trace behavioral responses to identity revelation over a meaningful time horizon and to rule out several alternative channels. Our study complements this literature by providing causal evidence on behavioral responses to sexual orientation disclosure in a naturally occurring virtual environment, leveraging a large-scale natural experiment, objective measures of revealed preferences, and the high degree of anonymity characteristic of online gaming platforms.
@@
The rest of the paper unfolds as follows. Section 2 describes the key elements of League of Legends that are relevant to our study and outlines the natural experiment we leverage to identify the effects of coming out. Section 3 introduces our data. Section 4 presents our main results. Section 5 examines the underlying mechanisms driving the estimated effects. Section 6 concludes.
@@ 2.1 League of Legends
League of Legends is a globally prominent multiplayer online game developed and published by Riot Games. Originally launched on October $27^{th}$, $2009$, the game attracted a vast player base, averaging $180$ million monthly active players as of $2022$ and peaking at $14$ million players in a single day. It has also achieved substantial financial success, generating $\mathdollar 1.8$ billion in revenue in that same year.111111 See, e.g., https://prioridata.com/data/league-of-legends.
League of Legends demands significant time and cognitive engagement from its user base. In our sample alone—which follows a subset of players over a seven-month period—users collectively spent 647,966647{,}966 hours in matches. Valuing this gameplay time at the 2022 U.S. federal minimum wage of $\mathdollar 7.25$ per hour implies an aggregate opportunity cost of $\mathdollar 4.7$ million. This figure represents a conservative lower bound of time spent on the game, as it excludes substantial time spent on pre- and post-match activities.
In League of Legends, players are divided into two teams of five players each to compete in matches with the aim of destroying the opposing team’s base. Players in each team coordinate to sort themselves into one of five roles, typically based on pre-selected role preferences and informal agreement in the pre-game virtual lobby. These roles represent crucial strategic positions, each requiring specific playstyles and contributing differently to the team’s final objective.
Before a match begins, players must select a playable character to control during the match.121212 The roster of playable characters in League of Legends expands over time as new characters are periodically released. At the time of our analysis, players could choose from $161$ characters. In our analysis, we measure players’ revealed preferences for a specific character by quantifying how frequently they select that character for their matches. Our objective is to investigate whether these preferences undergo any shifts following the disclosure of the character’s sexual orientation. Thus, we devote the rest of this section to exploring the design of characters in League of Legends and the process through which players select their characters for matches.
Each character has a unique set of skills and abilities and is specifically designed to excel in one or two of the distinct roles that players can take on within the team.131313 These roles are labeled top, jungle, mid, bottom, and support. Additionally, characters are crafted with a rich background that adds a narrative dimension to the game but does not have any impact on the game’s mechanics. This is achieved through the creation of detailed biographies and short stories that provide players with a deeper understanding of the character’s history and motivations, thus offering players the opportunity to connect with their chosen characters on a more personal level.141414 The list of all characters along with details about their abilities and histories is available at https://www.leagueoflegends.com/en-gb/champions/.